I am supposed to establish the trig identity csc x-sin x= cos x cot x!!! i do not get how to establish identities please help!!!!
Answer by Alan3354(69443)

csc x-sin x= cos x cot x
------------------------
Change it all to sines and cosines
---
(1/sin) - sin = cos*cos/sin
Multiply thru by sine
1 - sin^2 = cos^2
1 = sin^2 + cos^2 (Pythagorean Identity)
-------------------
The secret is to do something, to try.